OpenStudy (anonymous):

How can I solve this question Solve the equation S/R-3=T/R for R in terms of S and T

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

well start by gathering the r terms. so the equation becomes -3=t/r-s/r. But t/r-s/r=(t-s)/r so we have -3=(t-s)/r. Multiply both sides by r and then isolate r by dividing both sides by -3.